A 1, 500, 000 equipment has an estimated life of 20 years with a book value of 250, 000 at the end of the period. What is the depreciation charge and its book value after 10 years using the sinking fund method?
Answer by Theo(13342)

when you use the sinking fund method of depreciation, the annual interest rate for financing the asset needs to be used.
since you did not supply an interest rate, i will assume the rate is 10% per year for purposes of showing you how it works.

here's a reference on how the sinking method of depreciation works.
https://smallbusiness.chron.com/formulas-sinking-fund-method-depreciation-76525.html

the price of the asset is 1,500,000
the economic life of the asset is 20 years
the salvage value of the asset is 250,000 at the end of 20 years.

the present value of the asset is 1,500,000 minus the present value of 250,000 at 10% per year for 20 years.

the formula for the present value of the salvage is f = p * (1 + r) ^ n
solve for p to get:
p = f / (1 + r) ^ n
p is the present value
f is the future value r
r is the interest rate per time period
n is the number of time periods.

in this problem, the present value of the asset formula becomes:
p = 250,000 / 1.10 ^ 20 = 37,160.90701.

the present value of the asset is the price of the asset minus the present value of the salvage = 1,500,000 minus 37,160.90701 = 1,462,839.093.

the present value of the annuity formula is pv(annuity) = (1 - (1 + i) ^ -n) / i.
in this problem, that becomes pv(annuity) = (1 - 1.1^-n)/.1 = 8.51356372

the depreciation charge is the present value of the asset divided by the present value of the finance charge.
that is equal to 1,462,839.093 / 8.51356372 = 171,824.531.

that is also the payment at the end of each year for 20 years in a remaining balance account that starts with the asset price and ends with the salvage value.

the year by year transactions are shown in the following snapshot of the excel file worksheet that was used to calculate them.



instead of using the formula to calculate the depreciation charge, you can also use a financial calculator to determine what that charge should be.

a display of the use of that calculator is shown below:



in this calculator:

the present value is entered as -1500000.
the future value is entered as 250000.
the number of periods is entered as 20.
the interest rate % per period is entered as 10
the payment is made at the end of each period.

you click on pmt and the calculator tells you what the payment at the end of each year needs to be.
that's your depreciation charge.